# Fast sampling and model selection for Bayesian mixture models

### v1

@@ III.1 Algorithm 1
If $i$ was the only member of component $r$, delete component $r$ and decrease $k$ by 1.
Assemble a set of $k+1$ candidate states $\mu=(k,z)$, each of which has a weight $w_{\mu}$ associated with it, as follows. {enumerate*}
Of the $k+1$ candidate states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$. (We explicitly include the case $s=r$, where $i$ is placed back in the component $r$ it was just removed from, in which case there is no overall change of state.) The weights associated with each of these candidate states are
$w_{\mu}={N-k\over k}(n_{s}+1)P(k)P(x|k,z_{-i},z_{i}=s),$ (10)
for $s=1\ldots k$. Here $P(k)$ is the prior on $k$ as previously and $z_{-i}$ denotes all component assignments except for observation $i$, whose assignment is listed separately.
@@
where $k$ is the number of components before the new component is created.
@@
Once the set of target states is assembled, choose one of them $\mu$ with probability
$\pi_{\mu}={w_{\mu}\over\sum_{\nu}w_{\nu}},$ (12)
and update the system to that new state. Repeatedly performing these steps causes the distribution over states to converge from any initial state to the posterior distribution of the model—see Appendix A for details. An initial state that places all observations in a single component together is a simple choice that appears to work well. For common mixture models there are typically many factors that cancel from top and bottom of Eq. (12), giving a form that can be computed rapidly. We give an example in Section IV.1. Note that it is the likelihood that appears in the weights of Eqs. (10) and (11), and not, as one might imagine, the posterior probability. This is crucial to the correct working of the algorithm.
We give some results on the performance of this algorithm in Section IV.2. However, our primary interest in the algorithm is not for its own sake, but as a stepping stone to our second algorithm, which we now describe.
III.2 Algorithm 2
Our second algorithm builds on our first, adding some seemingly small changes that substantially improve performance. Given any initial choice of $k$ and assignment $z$ of the $N$ observations to components, algorithm 2 repeatedly performs the following steps. {enumerate*}
Remove $i$ from component $r$.
If $i$ was the only member of component $r$, delete component $r$ and decrease $k$ by 1.
Assemble a set of $k+1$ candidate states $\mu=(k,z)$, each of which has a weight $w_{\mu}$ associated with it, as follows. {enumerate*}
Of the $k+1$ states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$, as before. Now, however, the weights associated with these candidate states are
$w_{\mu}={N-k\over k}P(k)P(x|k,z_{-i},z_{i}=s).$ (13)
$w_{\mu}=kP(k+1)P(x|k+1,z_{-i},z_{i}=k+1),$ (14)
where $k$ is the number of components before the new component is created.
@@ III.2 Algorithm 2
Once the set of target states is assembled, choose one of them $\mu$ with probability $\pi_{\mu}$, Eq. (12), and update the system to that new state. A proof of the correctness of this algorithm is given in Appendix A. A crucial feature of the algorithm is that we choose a component at random in the first step—not an observation—then choose an observation from that component. This process, if repeated alone without the other parts of the algorithm, results in assignments drawn exactly from the distribution $P(z|k)$, Eq. (5), which is why the factor $n_{s}+1$ of Eq. (10) no longer appears in the weights. Moreover, these draws are “rejection-free,” meaning that every draw gives a new state, unlike the traditional Gibbs sampler, which often returns to the same state on successive samples—as much as 90% of the time in some cases. This change can significantly improve the efficiency of the algorithm. One way to understand this is to notice that the factor $n_{s}+1$ in (10) favors moving observations into large components, which means that an observation that starts out in a large component is more likely to be placed back in the same component again, resulting in a null move that does not update the state of the system. There are by definition more observations in a large component than a small one, so this inefficiency affects many observations on average. Our second algorithm does not suffer from this problem. The absence of the factor $n_{s}+1$ in Eq. (13) removes the bias toward large components and the uniform selection of a random component for each Monte Carlo move means that observations in small components are actually more likely to get picked than observations in large components. As we will see in Section IV.2, the result is a decrease in the number of null moves and a significantly faster mixing time. This behavior is, however, dependent on there being a range of component sizes in the first place. For data sets with components of roughly uniform size, the benefits of algorithm 2 will be more modest, and hence we expect the performance advantage to vary from one data set to another.
III.3 Parameter values
@@
with $P(x|k,z,\theta)$ as in Eq. (6). With $P(\theta|k)$ the appropriate canonical prior, Eq. (15) takes the same functional form as the prior, which generally makes it straightforward to estimate the expectation of $\theta$, its variance, modal value, and so forth. These estimates do still depend on $k,z$ and our Monte Carlo procedure returns many candidate values of $k,z$ and hence many values of $\theta$. For practical purposes one would often prefer just a single “best” value, which leads us to our next topic.
III.4 Component switching and
consensus components
@@ III.4 Component switching and
An issue with Monte Carlo sampling from mixture models is that the component labels themselves are meaningless. If we take a component assignment $z$ and permute the $k$ component labels we still have the same division of the observations, and hence all permutations have the same posterior probability. This means for instance that the marginal probability $P(z_{i}=r|x,k)$ of observation $i$ belonging to component $r$ is always a constant $1/k$ independent of the data $x$, so one cannot meaningfully ask “What component does $i$ belong to?” The components are not identifiable. This problem shows up in the Monte Carlo results as “component switching”—the algorithm may sample the same or similar component assignments but with different permutations of the labels, so that the similarity is difficult to see.
@@
The method has not failed. It is drawing correctly from the posterior distribution, but for practical purposes it is not giving us type of answer we are looking for. In most cases the practitioner would prefer a single unambiguous assignment of observations to components, or at least a set of closely similar ones. Methods that employ EM algorithms deal with this by selecting a single estimate of the parameters $\theta$ and then reporting component assignments conditioned on those values. This breaks the symmetry over permutations, but at the expense of restricting us to point estimates of the parameters, which has its own shortcomings, as discussed in the introduction.
@@
What can we do instead? Some investigators have suggested imposing symmetry-breaking rules on the component assignments so that only one assignment in each orbit of the permutation group is sampled RG97 ; Stephens00 , or choosing the permutations that make the sampled assignments most similar ZDA23 . These approaches work well when the posterior distribution is concentrated around the modal component assignment, but fail when the distribution is broad enough that its faithful representation would require frequent component switching. In the latter situation, algorithms that prohibit component switching give biased estimates.
@@
A better approach, in our opinion, is to sample assignments without restriction, allowing component switching, and then, if a single assignment is desired, to construct one after the fact as the consensus clustering that best captures the common features of all the samples. Many methods for doing this have been proposed MPMG03 ; Bryant03 ; GF08 ; VR11 ; LF12 ; ZDA23 and we give some example applications in Section IV.4. However, it is important to understand that it may not be possible to capture all the salient features of the sampled assignments in a single consensus configuration—there may be multiple configurations that have high posterior probability but which differ significantly. This kind of behavior is observed for instance in network community detection GDC10 , and has led to the development of alternative approaches, such as “building block” decompositions RN20 ; Arthur24 or methods for finding multiple representative assignments within a large sample KN22 .
@@
Alternatively, we can simply restrict ourselves to analyses based purely on quantities that are invariant under permutations of the component labels. Examples of such quantities include the coincidence rate—the frequency with which two observations belong to the same component—and the mutual information between sampled component assignments for two observations, or between the component assignments and the data. There are problems of interest that can be tackled directly using such quantities. An example is variable selection, the problem of deciding which observed variables are most informative about component membership. We give an example in Section IV.4.
IV Results
In this section we give results on the performance and applications of our methods. Our algorithms can be applied to a range of mixture models, but for purposes of illustration we here focus on the example of latent class analysis.
IV.1 Latent class analysis

### v2

@@ 3 Monte Carlo algorithm
  3. 3.
    If $i$ was the only member of component $r$, delete component $r$, relabel component $k$ to be the new component $r$, and decrease $k$ by 1. (If $r=k$ then no relabeling is necessary; we simply delete the component.)
  4. 4.
    Assemble a set of $k+1$ candidate states, each of which has a weight associated with it, as follows.
    1. (a)
      Of the $k+1$ states, $k$ of them are states in which $i$ is placed in one of the $k$ current components $s$. (We explicitly include the case $s=r$, where $i$ is placed back in the component $r$ it was just removed from, in which case there is no overall change of state.) The weights associated with each of these candidate states are
      $w_{s}={P(x|k,z_{-i},z_{i}=s)\over P(x|k,z_{-i})},$ (14)
      where $z_{-i}$ denotes the component assignments for all observations except for observation $i$.
    2. (b)
@@
      where $P(k)$ is the prior on $k$ as previously, and $k$ is the number of components before the new component is added.
  5. 5.
@@
    Once the set of target states is assembled, choose one of them with probability
    $p_{s}={w_{s}\over\sum_{s}w_{s}}$ (16)
    and update the system to that new state.
@@
A proof of the correctness of this algorithm is given in Appendix A. A crucial feature of the algorithm is that we choose a component in the first step, not an observation, then choose an observation from that component. This process, if repeated alone without the other parts of the algorithm, results in assignments drawn exactly from the prior $P(z|k)$, Eq. (9), which is why no term for the prior appears in the formulas for the weights. Moreover, these draws are “rejection-free,” meaning that no draws are discarded, as they are for instance in a Metropolis-Hastings process. This change significantly improves the efficiency of the algorithm. One way to understand the improvement is to notice that, if the prior were included in the weights, the factor of $n_{r}!$ in Eq. (9) would favor moving observations into large components, which would mean that an observation that started out in a large component would be more likely to be placed back in the same component again, resulting in a wasted move that did not update the state of the system. Furthermore, the uniform selection of an observation at the start of each Monte Carlo move, as in a conventional Gibbs sampler, would result in a bias toward large components that would emphasize such wasted moves over moves that actually generate new states. This issue has been noted by previous authors, such as Miller and Harrison [37], who suggest the split-merge sampling method of Jain and Neal [26] as a potential remedy. Our algorithm, however, offers a simpler solution. By moving the prior out of the weights and into the proposal mechanism, Monte Carlo moves that place observations in large components are no longer favored and the uniform selection of a random component at the start of each move eliminates the selection bias in favor of large components.
As a specific example where this approach can make a difference, consider what happens when the algorithm creates a new component, which initially is small by definition. Creation of components is a natural part of the operation of any algorithm that samples from the posterior distribution of the number of components, but inevitably the algorithm will sometimes create components where they are not justified by the data, in which case they will normally be immediately deleted again. With a traditional Gibbs sampler (or collapsed Gibbs sampler), the process of deletion is slow because an algorithm that updates all observations with equal frequency updates the observations in a small component of size $\mathrm{O}(1)$ on only a fraction $\mathrm{O}(1/N)$ of Monte Carlo steps. Our algorithm, on the other hand, since it chooses a random component on every step, performs such updates $\mathrm{O}(1/k)$ of the time, and hence can be expected to remove unwanted components a factor of $\mathrm{O}(N/k)$ faster than the traditional algorithm. For large data sets this factor can be on the order of thousands or more.
It would be possible to speed up our algorithm further by parallelization. The bulk of the computational effort goes into calculating the weights in Eqs. (14) and (15), which has to be done anew for every Monte Carlo step. The individual weights however are not dependent and can be calculated in any order, which makes the algorithm trivially parallelizable and an ideal candidate for a multithreaded or vectorized implementation.
3.1 Parameter values
@@
with $P(x|k,z,\theta)$ as in Eq. (10). If $P(\theta|k)$ is the appropriate canonical prior, Eq. (17) takes the same functional form as the prior, which generally makes it straightforward to estimate the expectation of $\theta$, its variance, modal value, and so forth. These estimates do still depend on $k,z$ and our Monte Carlo procedure returns many candidate values of $k,z$ and hence many values of $\theta$. For practical purposes one would often prefer just a single “best” value, which leads us to our next topic.
3.2 Label switching and consensus components
@@
An issue with sampling from mixture models is that the component labels themselves are meaningless. If we take a component assignment $z$ and permute the $k$ component labels we still have the same division of observations, and hence all permutations have the same posterior probability. This means for instance that the marginal probability $P(z_{i}=r|x,k)$ of observation $i$ belonging to component $r$ is always a constant $1/k$ independent of the data $x$, so one cannot meaningfully ask “What component does $i$ belong to?” The components are not identifiable. This problem is manifested in the Monte Carlo results as “component switching” or “label switching”—the algorithm may sample the same or similar component assignments but with different permutations of the labels, so that the similarity is difficult to see.
@@
The method has not failed. It is drawing correctly from the posterior distribution, but for practical purposes it is not giving us the type of answer we are looking for. In most cases the practitioner would prefer a single unambiguous assignment of observations to components, or at least a set of closely similar ones. Methods that employ EM algorithms deal with this by selecting a single estimate of the parameters $\theta$ and then reporting component assignments conditioned on those values. This breaks the symmetry over permutations, but the EM algorithm has other shortcomings, as discussed in the introduction.
@@
What can we do instead? Some investigators have suggested imposing symmetry-breaking rules on the component assignments so that only one assignment in each orbit of the permutation group is sampled [50, 55], or choosing the permutations that make the sampled assignments most similar [63]. These approaches work well when the posterior distribution is concentrated around the modal component assignment, but fail when the distribution is broad enough that its faithful representation would require frequent label switching. In the latter situation, algorithms that prohibit label switching give biased estimates [17].
@@
A better approach, in our opinion, is to sample assignments without restriction, allowing label switching, and then, if a single assignment is desired, to construct one after the fact as the consensus clustering that best captures the common features of all the samples. Many methods for doing this have been proposed [38, 8, 19, 59, 29, 63] and we give an example application in Section 4.3. It is worth noting, however, that it may not be possible to capture all the salient features of the sampled assignments in a single consensus configuration—there may be multiple configurations that have high posterior probability but which differ significantly. This kind of behavior has been observed for instance in network community detection [20], and has led to the development of alternative approaches, such as “building block” decompositions [51, 1] or methods for finding multiple representative assignments within a large sample [28].
@@
Alternatively, we can simply restrict ourselves to analyses based purely on quantities that are invariant under permutations of the component labels. Examples of such quantities include the coincidence rate—the frequency with which two observations belong to the same component—and the mutual information between sampled component assignments for two observations, or between the component assignments and the data. There are problems of interest that can be tackled directly using such quantities. An example is variable selection, the problem of deciding which observed variables are most informative about component membership. We give an example in Section 4.3.
4 Results
In this section we give results on applications and performance of our methods for a variety of models.
4.1 Gaussian mixtures
For our first example, we perform benchmark tests using perhaps the simplest of possible models, a univariate Gaussian mixture with constant variance. This model is fully parametrized by the means $\mu_{1}\ldots\mu_{k}$ of the components, plus the number of components $k$ and assignments $z$. The data likelihood is
$P(x|k,z,\mu)$ $=\prod_{i}{1\over\sqrt{2\pi\sigma^{2}}}\exp\biggl[-{(x_{i}-\mu_{z_{i}})^{2}\over 2\sigma^{2}}\biggr]$
$=(2\pi\sigma^{2})^{-N/2}\prod_{r}\exp\biggl[-{\sum_{i\in r}(x_{i}-\mu_{r})^{2}\over 2\sigma^{2}}\biggr]$
$=(2\pi\sigma^{2})^{-N/2}\prod_{r}\exp\biggl[-{n_{r}\over 2\sigma^{2}}\bigl[(\langle x\rangle_{r}-\mu_{r})^{2}+\sigma_{r}^{2}\bigr]\biggr],$ (18)
where $\langle x\rangle_{r}$ and $\sigma_{r}^{2}$ denote the mean and variance of the observations within component $r$, and $n_{r}$ is the number of observations in component $r$ as previously. Integrating with respect to $\mu_{r}$ over an interval of width $a$, with $a\gg 2\pi\sigma^{2}$ and a uniform prior, the marginal likelihood is then
$P(x|k,z)={1\over a^{k}(2\pi\sigma^{2})^{(N-k)/2}}\prod_{r}{1\over\sqrt{n_{r}}}\,\mathrm{e}^{-n_{r}\sigma_{r}^{2}/2\sigma^{2}}.$ (19)
For the prior $P(k)$ on the number of classes we make the uniform choice $P(k)=1/N$, then the weight of Eq. (14) for Monte Carlo moves that do not create a new component becomes
$w_{s}=\sqrt{n_{s}/n_{s}^{\prime}\over 2\pi\sigma^{2}}\,\mathrm{e}^{-(n_{s}^{\prime}{\sigma_{s}^{\prime}}^{2}-n_{s}\sigma_{s}^{2})/2\sigma^{2}}.$ (20)
With a little extra work we can show that
$n_{s}^{\prime}{\sigma_{s}^{\prime}}^{2}-n_{s}\sigma_{s}^{2}={n_{s}\over n_{s}^{\prime}}(x_{i}-\langle x\rangle_{s})^{2},$ (21)
and putting $n_{s}^{\prime}=n_{s}+1$, we have
$w_{s}=\sqrt{n_{s}/(n_{s}+1)\over 2\pi\sigma^{2}}\,\exp\biggl[-{n_{s}/(n_{s}+1)\over 2\sigma^{2}}(x_{i}-\langle x\rangle_{s})^{2}\biggr].$ (22)
Performing the same series of steps for Eq. (15), the weight for the remaining move which creates a new component $k+1$ is simply
$w_{k+1}={k^{2}/a\over N-k}.$ (23)
This simple model provides a convenient arena for benchmarking the performance of our algorithm against competing methods. For these tests we generate synthetic data sets of N=10  [^0] N=10\,000 observations each, divided equally between $k$ Gaussian components with the observations in component $r=1\ldots k$ drawn from $x_{i}\sim\text{Normal}(3k,1)$. The performance of Monte Carlo algorithms for this problem typically varies with $k$, either because of the time taken for individual Monte Carlo steps or because of varying mixing times, or both, so we report results for a range of $k$ values to give a complete picture.
We compare the performance of our algorithm against three other popular Monte Carlo methods for Bayesian mixture models, as implemented in the software package BayesMix [4]: Neal’s second and third algorithms, which are traditional Gibbs samplers [41] and the split-merge algorithm of Jain and Neal [26]. Results are given in Table 1.
Table 1: Running time per sweep in milliseconds and estimated mixing time in sweeps for the algorithm of this paper and a selection of competing algorithms, applied to a Gaussian model with synthetic data and various numbers of components $k$. Figures in parentheses indicate standard errors on the trailing digits.
This paper Neal’s algorithm 2 Neal’s algorithm 3 Split-merge
$k$ Time Mixing Time Mixing Time Mixing Time Mixing
3 1.4 21.2(8.4) 17.9 120.4(11.7) 23.6 144.9(16.7) 139.3 18.7(2.5)
5 2.2 25.4(6.2) 27.0 79.8(10.3) 23.8 84.2(11.7) 131.2 23.9(2.2)
7 3.7 20.8(3.3) 29.2 80.9(7.9) 23.9 60.4(3.5) 129.6 11.2(1.3)
10 4.5 24.0(2.7) 34.1 49.9(3.4) 22.5 43.5(5.1) 152.8 8.1(5)
BayesMix is written in the C++ language for speed and is a well optimized piece of software, but it can only be as efficient as the algorithms it implements, and in this respect it is laboring under two disadvantages. First, the algorithms are complex. Even a single Monte Carlo step can require a large amount of computation. For instance, the fastest of the competing algorithms we consider is Neal’s second algorithm. In our tests we measure progress in “sweeps,” meaning $N$ Monte Carlo updates of individual observations, so that each observation is updated once on average per sweep. For the case where $k=5$, a complete sweep of Neal’s second algorithm takes 27.0 milliseconds of CPU time on the hardware used for these tests. For the algorithm of this paper, by contrast, a complete sweep takes 2.2 milliseconds on the same hardware, less than a tenth of the time. Not all of this difference will be due to the complexity of the algorithm. Some is due to complexity of the models themselves or to overheads inherent in the architecture of the BayesMix package, although the authors state that such overheads are small. Nonetheless, for a practitioner applying these methods the differences are real and will have a substantial impact on running times.
The second disadvantage of these algorithms is that their mixing time can be slow, meaning it takes many successive Monte Carlo sweeps to generate one statistically independent sample of the system. The number of sweeps needed to generate an independent sample can be quantified by calculating the correlation time of the sample chain. In this paper we compute the autocorrelation of the log-likelihood and then from it calculate the integrated correlation time [42]. Figures for the resulting times, measured in sweeps and averaged over ten repetitions with randomly generated data, are given in Table 1. Here again we see significant differences between algorithms. Once more taking the example of Neal’s second algorithm for the case $k=5$, the correlation time is measured to be about 80 Monte Carlo sweeps, while for the algorithm of this paper it is about 25 sweeps, less than a third as much, meaning that it takes significantly fewer steps to generate one independent sample configuration. The combination of faster sweeps and faster mixing makes the algorithm of this paper about 38 times faster overall than Neal’s algorithm in this particular test, enough to make the difference between a calculation that runs in an hour and one that takes just a minute or two. A difference of this magnitude is not merely a quantitative one, but a qualitative one also: the ability to obtain useful results in minutes or seconds makes possible forms of interactive and exploratory data analysis that are impractical with slower methods.
For the other algorithms the details vary but the overall outcome is similar. Consider, for instance, the split-merge algorithm. This powerful algorithm uses nonlocal moves that split and merge whole components to achieve in a single Monte Carlo step what in other algorithms takes many steps. This improves mixing time, but at the expense of greater computational effort. Looking again at the case of $k=5$, the correlation time for the split-merge algorithm is significantly better than Neal’s algorithm—only 24 sweeps, which is also slightly better than our own algorithm at 25 sweeps. On the other hand, each sweep of the system with N=10  [^0] N=10\,000 observations takes 131 milliseconds for the split-merge algorithm, compared with just 2.2 milliseconds for our algorithm, so that our algorithm is still faster overall by a factor of about 56.
4.2 Poisson mixtures
For our next example we consider a mixture of Poisson distributions over integer observations $x_{i}$, such that the data likelihood is
$P(x|k,z,\mu)=\prod_{i}{\mu_{z_{i}}^{x_{i}}\over x_{i}!}\exp(-\mu_{z_{i}})\\ =\biggl[\prod_{i}{1\over x_{i}!}\biggr]\biggl[\prod_{r=1}^{k}\mu_{r}^{\sum_{i\in r}x_{i}}\mathrm{e}^{-n_{r}\mu_{r}}\biggr].$
Again the model is completely parametrized by the means $\mu_{r}$ of the components, along with the number $k$ of components and the assignments $z_{i}$. Assuming a conventional gamma prior on the means of the form
$P(\mu)={\beta^{\alpha}\over\Gamma(\alpha)}\mu^{\alpha-1}\mathrm{e}^{-\beta\mu}$ (24)
and integrating, we find the marginal likelihood to be
$P(x|k,z)=\prod_{i}{1\over x_{i}!}\prod_{r=1}^{k}{\beta^{\alpha}\Gamma(\sum_{i\in r}x_{i}+\alpha)\over\Gamma(\alpha)(n_{r}+\beta)^{\sum_{i\in r}x_{i}+\alpha}}.$ (25)
Substituting this expression into Eqs. (14) and (15), we get Monte Carlo weights
$w_{s}={\Gamma(X_{s}+x_{i}+\alpha)\over\Gamma(X_{s}+\alpha)}\,{(n_{s}+\beta)^{X_{s}+\alpha}\over(n_{s}+\beta+1)^{X_{s}+x_{i}+\alpha}},\qquad X_{s}=\sum_{i\in s}x_{i}$ (26)
for steps that move an observation $i$ to an existing component $s$, and
$w_{k+1}={k^{2}\over N-k}\,{P(k+1)\over P(k)}\,{\Gamma(x_{i}+\alpha)\over\Gamma(\alpha)}\,{\beta^{\alpha}\over(\beta+1)^{x_{i}+\alpha}}$ (27)
for steps that move observation $i$ to newly created component $k+1$.
As an example, Fig. 1 shows results for a clinical data set from Wang et al. [61] recording the number of seizures experienced per day by an epileptic patient under varying treatment regimes over a 140-day observation period. Wang et al. analyzed the data using two- and three-component Poisson mixture models. For our analysis we applied the Monte Carlo algorithm above with $\alpha=1$ and $\beta=0.01$ for 10 000 sweeps of burn-in followed by 100 000 sweeps for sampling, the entire calculation taking about three seconds on the author’s laptop. Figure 1 shows the estimated posterior distribution over the number of components, which is simply a histogram of the sampled values of $k$ and, as the figure reveals, fits with both two and three components are firmly ruled out: out of 100 000 samples, zero of them had either $k=2$ or $k=3$. The smallest observed number of components was $k=4$ and the smallest plausible value (at a $p=0.05$ significance level) was $k=5$. The modal value was $k=7$. Realistically, this implies one of two things: either the original authors’ assumption of two or three components was wrong or the daily seizure events are not Poisson distributed and a mixture of Poissons is a poor model.
Figure 1: Estimated distribution of the number of components in a fit of the epileptic seizures data set of Wang et al. [61] to the Poisson mixture model described in the text.
Figure 2 shows another, more whimsical application, to a data set that comes from a candy dispenser in the hallway outside the author’s office—a machine that dispenses candy at the push of a button. Allegations have swirled for some time that this machine is not fair: sometimes (it is alleged) it dispenses a generous handful of candies, sometimes a miserly pittance. The data set represents real counts of the number of candies dispensed on 853 pushes of the button, which were analyzed using the Poisson model above with 10 000 sweeps of burn-in followed by 100 000 sweeps of sampling, which takes about six seconds of running time. The figure shows the distribution of numbers of candies (main figure) and the inferred number of components (inset). As the inset shows, the results firmly exclude the possibility that there is only one component and lean heavily into the hypothesis that there are two: a miserly one (mean 7.382(2), 54.7% of observations) and a generous one (mean 15.987(3), 45.3% of observations). The main figure shows the observed distribution of numbers of candies along with the fitted distribution inferred from the model. Together these results strongly support the charge that the machine is inequitable in the performance of its duties.
Figure 2: Main figure: The empirical distribution (histogram) and fitted distribution (points) for the number of candies dispensed by the machine. Inset: The posterior distribution of the number of components estimated from the Monte Carlo results.
4.3 Latent class analysis